# EvoIR: Towards All-in-One Image Restoration via Evolutionary Frequency Modulation

### v1

@@ Abstract
All-in-One Image Restoration (AiOIR) tasks often involve diverse degradation that require robust and versatile strategies. However, most existing approaches typically lack explicit frequency modeling and rely on fixed or heuristic optimization schedules, which limit the generalization across heterogeneous degradation. To address these limitations, we propose EvoIR, an AiOIR‑specific framework that introduces evolutionary frequency modulation for dynamic and adaptive image restoration. Specifically, EvoIR employs the Frequency-Modulated Module (FMM) that decomposes features into high- and low-frequency branches in an explicit manner and adaptively modulates them to enhance both structural fidelity and fine-grained details. Central to EvoIR, an Evolutionary Optimization Strategy (EOS) iteratively adjusts frequency-aware objectives through a population-based evolutionary process, dynamically balancing structural accuracy and perceptual fidelity. Its evolutionary guidance further mitigates gradient conflicts across degradation and accelerates convergence. By synergizing FMM and EOS, EvoIR yields greater improvements than using either component alone, underscoring their complementary roles. Extensive experiments on multiple benchmarks demonstrate that EvoIR outperforms state-of-the-art AiOIR methods.
@@ 1 Introduction
Image restoration recovers a high-quality image from its degraded observation. Traditionally, this problem has been tackled by task-specific networks, each tailored to a particular type of degradation. Such task-specific models have demonstrated impressive performance across various tasks, including denoising Shen et al. (2023), dehazing Song et al. (2023), deraining Chen et al. (2023a), deblurring Tsai et al. (2022), low-light enhancement Ma et al. (2023b), and under-water enhancement Zhang et al. (2025b).
However, task-specific methods suffer from limited generalization, as they are inherently tailored to handle only predefined degradation types. When applied to unfamiliar degradation, their performance tends to degrade dramatically. General image restoration approaches have been proposed Zamir et al. (2021); Chen et al. (2022); Cui et al. (2023a); Xia et al. (2023) to address these limitations. Although these models are capable of addressing various degradation types, they are generally trained and tested on single tasks, which limits their practicality in real-world settings that involve complex and mixed degradation.
Recently, All-in-One image restoration methods Ai et al. (2024); Conde et al. (2024); Liu et al. (2025); Cui et al. (2025); Zamfir et al. (2025); Tian et al. (2025) have emerged as promising solutions to the limitations aforementioned. These approaches restore images corrupted by multiple degradation types within a unified framework. Early efforts such as AirNet Li et al. (2022) constructed explicit degradation encoders to obtain discriminative degradation-aware features. Subsequent works, including ProRes Ma et al. (2023a) and PromptIR Potlapalli et al. (2023), enhanced performance by incorporating visual prompts as guidance. The work in Tan et al. (2024) exploits the rich feature representations of large-scale vision models, such as CLIP Radford et al. (2021) and DINO Caron et al. (2021). Perceive-IR Zhang et al. (2025a) formulates image restoration from a quality-aware perspective, enabling the model to adjust its restoration strategy based on degradation severity.
@@ 2.1 All-in-One Image Restoration
All-in-One image restoration methods address diverse degradation using a unified model, offering improved storage and deployment efficiency over task-specific Yasarla & Patel (2019); Chen et al. (2023a); Kupyn et al. (2018); Cai et al. (2023); Wu et al. (2022) and general-purpose Chen et al. (2022); Zamir et al. (2022b); Xia et al. (2023); Guo et al. (2024) approaches.
The core challenge lies in restoring multiple degradation types within a shared parameter space. To address this, AirNet Li et al. (2022) applies contrastive learning for degradation discrimination, IDR Zhang et al. (2023) adopts a two-stage ingredient-oriented design, and ProRes Ma et al. (2023a) introduce visual prompts for guided restoration. Recent methods Tan et al. (2024); Luo et al. (2023) further leverage large-scale pre-trained vision models to enhance texture and semantics.
@@ 2.2 Frequency Domain-Based Image Restoration
Recent studies have emphasized the importance of frequency-domain modeling in boosting restoration performance Cui et al. (2023b); Wu et al. (2025). CSNet Cui et al. (2024) integrates channel-wise Fourier transforms and multi-scale spatial frequency modules, guided by frequency-aware loss, to enhance both spectral interaction and spatial detail. For deblurring, the Efficient Frequency Domain Transformer Kong et al. (2023) reformulates attention and feed-forward layers in the frequency domain, improving visual quality and efficiency. FPro Zhou et al. (2024) employs frequency decomposition and prompt learning to guide structure–detail recovery across diverse tasks. AdaIR Cui et al. (2025) further mines degradation-specific frequency priors and applies bidirectional modulation to enhance reconstruction.
@@ 3.1 Overall Pipeline
Backbone architecture. Following prior frequency‑domain designs Kong et al. (2023), we extend them into the RES‑FFTB module with multi‑head self‑attention and residual connections for efficient information flow. We also adopt the Refine Block from AdaIR Cui et al. (2025) to further polish the fused features before decoding.
@@ 4.1 Experimental Setup
Datasets: Following Li et al. (2022); Potlapalli et al. (2023), we consider One-by-One training paradigm (single-task training), All-in-One training paradigm (multi-task joint training), composited degradation and remote sensing imagery. For One-by-One and All-in-One, we explore two common degradation combinations: N+H+R (Noise, Haze, Rain) and N+H+R+B+L (Noise, Haze, Rain, Blur, Low-light). Denoising: BSD400 Arbelaez et al. (2010), WED Ma et al. (2016), CBSD68 Martin et al. (2001) and Kodak24 Franzen (1999); Dehazing: OTS from RESIDE-$\beta$ Li et al. (2018) and SOTS-Outdoor Li et al. (2018); Deraining: Rain100L Yang et al. (2017); Deblurring: GoPro Nah et al. (2017); Low-light Enhancement: LOL Wei et al. (2018). Details of these datasets are summarized in appendix. For composited degradation, we use CDD11 dataset DBLP:conf/eccv/GuoGLZLH24. For remote sensing AiOIR, we choose MDRS-LandsatDBLP:journals/inffus/LiheYHJXCSZ25. For these two settings, we follow the original split of training and test sets.
@@
Following Cui et al. (2025), we adopt task-specific resampling ratios to address data imbalance across restoration tasks. Specifically, the data expansion ratios of 3, 120, 5, and 200 are applied to denoising, deraining, deblurring, and low-light enhancement, respectively, while dehazing remains unaltered. Training is conducted on 4 NVIDIA A100 40GB GPUs using cropped patches of size $128\times 128$, with random horizontal and vertical flipping for data augmentation.
Table 1: Performance comparison (PSNR/SSIM) under All-in-One (“N+H+R”) setting. Results are partially sourced from Perceive-IR Zhang et al. (2025a).
@@ 4.2 All-in-One Restoration Results
Tab.1 shows that EvoIR achieves the best average PSNR (33.00 dB) and SSIM (0.922). Compared to the three latest AiOIR methods: AdaIR Cui et al. (2025), MoCE-IR Zamfir et al. (2025) and DFPIR Tian et al. (2025), our method improves the average PSNR/SSIM by +0.31 dB/+0.004, +0.23 dB/+0.005 and +0.12 dB/+0.003, respectively. Notably, EvoIR achieves new state-of-the-art results on dehazing (32.08/0.982) and deraining (39.07/0.985), demonstrating the effectiveness of adaptive frequency modulation and evolutionary optimization.
@@
Table 2: Performance comparison (PSNR/SSIM) under All-in-One (“N+H+R+B+L”) setting. Denoising results only reports noise level under $\sigma=25$ following Zhang et al. (2023). Results are partially sourced from Perceive-IR Zhang et al. (2025a).
@@ Appendix B Code and Models
Code and models for evaluation can be found here.
@@ C.2 Baselines
All-in-One setting: For “N+H+R” setting, we include the following recent All-in-One models: AirNet Li et al. (2022), IDR Zhang et al. (2023), ProRes Ma et al. (2023a), PromptIR Potlapalli et al. (2023), NDR Yao et al. (2024), Gridformer Wang et al. (2024), InstructIR Conde et al. (2024), Up-Restorer Liu et al. (2025), Perceive-IR Zhang et al. (2025a), AdaIR Cui et al. (2025), MoCE-IR Zamfir et al. (2025), DFPIR Tian et al. (2025). For “N+H+R+B+L” setting, we also include TAPE Liu et al. (2022), TransWeather Valanarasu et al. (2022).
One-by-One setting: We adopt task-specific and general methods tailored to individual degradation types. For Denoising: DnCNN Zhang et al. (2017), FFDNet Zhang et al. (2018), ADFNet Shen et al. (2023), MIRNet-v2 Zamir et al. (2022a), Restormer Zamir et al. (2022b), NAFNet Chen et al. (2022); Dehazing: adding DehazeNet Cai et al. (2016), FDGAN Dong et al. (2020), DehazeFormer Song et al. (2023), FSNet Cui et al. (2023a); Deraining: adding UMR Yasarla & Patel (2019), LPNet Gao et al. (2019), DRSformer Chen et al. (2023a); Deblurring: adding DeblurGAN Kupyn et al. (2018), Stripformer Tsai et al. (2022), HI-Diff Chen et al. (2023b), MPRNet Zamir et al. (2021); Low-light enhancement: adding URetinex Wu et al. (2022), SMG Xu et al. (2023), Retinexformer Cai et al. (2023), MIRNet Zamir et al. (2020), DiffIR Xia et al. (2023). We report All-in-One methods retrained under the One-by-One setting for comparison.
@@ E.2 Visual Comparison under Three Degradations
Due to page limitations, we provide additional visual comparisons for the three-degradation setting (“N+H+R”). As shown in Fig. 6, 7, 8, 9, and 10, we compare our EvoIR against recent state-of-the-art methods, including AdaIR Cui et al. (2025) (ICLR’25), DFPIR Tian et al. (2025) (CVPR’25), and MoCE-IR Zamfir et al. (2025) (CVPR’25).
@@ E.3 Visual Comparison under Five Degradations
As some methods do not provide visual results under the five-degradation setting (“N+H+R+B+L”), we select InstructIR Conde et al. (2024) (ECCV’24) and MoCE-IR Zamfir et al. (2025) (CVPR’25) for comparison with our EvoIR.

### v2

@@ Abstract
All-in-One Image Restoration (AiOIR) tasks often involve diverse degradation that require robust and versatile strategies. However, most existing approaches typically lack explicit frequency modeling and rely on fixed or heuristic optimization schedules, which limit the generalization across heterogeneous degradation. To address these limitations, we propose EvoIR, an AiOIR-specific framework that introduces evolutionary frequency modulation for dynamic and adaptive image restoration. Specifically, EvoIR employs the Frequency-Modulated Module (FMM) that decomposes features into high- and low-frequency branches in an explicit manner and adaptively modulates them to enhance both structural fidelity and fine-grained details. Central to EvoIR, an Evolutionary Optimization Strategy (EOS) iteratively adjusts frequency-aware objectives through a population-based evolutionary process, dynamically balancing structural accuracy and perceptual fidelity. Its evolutionary guidance further mitigates gradient conflicts across degradation and accelerates convergence. By synergizing FMM and EOS, EvoIR yields greater improvements than using either component alone, underscoring their complementary roles. Extensive experiments on multiple benchmarks demonstrate that EvoIR outperforms state-of-the-art AiOIR methods. Code and models for evaluation are released: https://github.com/leonmakise/EvoIR.
@@ 1 Introduction
Image restoration recovers a high-quality image from its degraded observation. Traditionally, this problem has been tackled by task-specific networks, each tailored to a particular type of degradation. Such task-specific models have demonstrated impressive performance across various tasks, including denoising ADFNet, dehazing DehazeFormer, deraining DRSformer, deblurring Stripformer, low-light enhancement DBLP:journals/nn/MaWZZ23, and under-water enhancement DBLP:journals/corr/abs-2501-12981.
However, task-specific methods suffer from limited generalization, as they are inherently tailored to handle only predefined degradation types. When applied to unfamiliar degradation, their performance tends to degrade dramatically. General image restoration approaches have been proposed MPRNet; NAFNet; FSNet; DiffIR to address these limitations. Although these models are capable of addressing various degradation types, they are generally trained and tested on single tasks, which limits their practicality in real-world settings that involve complex and mixed degradation.
Recently, All-in-One image restoration methods DBLP:conf/cvpr/AiHZW024; InstructIR; DBLP:conf/aaai/LiuYL025; DBLP:conf/iclr/0001ZKKSK25; DBLP:conf/cvpr/ZamfirWMTP0T25; DBLP:conf/cvpr/TianLLLR25 have emerged as promising solutions to the limitations aforementioned. These approaches restore images corrupted by multiple degradation types within a unified framework. Early efforts such as AirNet AirNet constructed explicit degradation encoders to obtain discriminative degradation-aware features. Subsequent works, including ProRes ProRes and PromptIR PromptIR, enhanced performance by incorporating visual prompts as guidance. The work in CLIP_AWR exploits the rich feature representations of large-scale vision models, such as CLIP CLIP and DINO DINO. Perceive-IR DBLP:journals/corr/abs-2408-15994 formulates image restoration from a quality-aware perspective, enabling the model to adjust its restoration strategy based on degradation severity.
@@ 2.1 All-in-One Image Restoration
All-in-One image restoration methods address diverse degradation using a unified model, offering improved storage and deployment efficiency over task-specific UMR; DRSformer; DeblurGAN; Retinexformer; URetinex and general-purpose NAFNet; Restormer; DiffIR; MambaIR approaches.
The core challenge lies in restoring multiple degradation types within a shared parameter space. To address this, AirNet AirNet applies contrastive learning for degradation discrimination, IDR IDR adopts a two-stage ingredient-oriented design, and ProRes ProRes introduce visual prompts for guided restoration. Recent methods CLIP_AWR; DA-CLIP further leverage large-scale pre-trained vision models to enhance texture and semantics.
@@ 2.2 Frequency Domain-Based Image Restoration
Recent studies have emphasized the importance of frequency-domain modeling in boosting restoration performance DBLP:conf/iclr/0001TBRGC0K23; DBLP:journals/pr/WuLWLH25. CSNet DBLP:conf/ijcai/0001LRK24 integrates channel-wise Fourier transforms and multi-scale spatial frequency modules, guided by frequency-aware loss, to enhance both spectral interaction and spatial detail. For deblurring, the Efficient Frequency Domain Transformer DBLP:conf/cvpr/KongDGLP23 reformulates attention and feed-forward layers in the frequency domain, improving visual quality and efficiency. FPro DBLP:conf/eccv/ZhouPSCQY24 employs frequency decomposition and prompt learning to guide structure–detail recovery across diverse tasks. AdaIR DBLP:conf/iclr/0001ZKKSK25 further mines degradation-specific frequency priors and applies bidirectional modulation to enhance reconstruction.
@@ 3.1 Overall Pipeline
Backbone architecture. Following prior frequency‑domain designs DBLP:conf/cvpr/KongDGLP23, we extend them into the RES‑FFTB module with multi‑head self‑attention and residual connections for efficient information flow. We also adopt the Refine Block from AdaIR DBLP:conf/iclr/0001ZKKSK25 to further polish the fused features before decoding.
@@ 4.1 Experimental Setup
Datasets: Following AirNet; PromptIR, we consider One-by-One training paradigm (single-task training), All-in-One training paradigm (multi-task joint training), composited degradation and remote sensing imagery. For One-by-One and All-in-One, we explore two common degradation combinations: N+H+R (Noise, Haze, Rain) and N+H+R+B+L (Noise, Haze, Rain, Blur, Low-light). Denoising: BSD400 BSD400, WED WED, CBSD68 BSD68 and Kodak24 Kodak24; Dehazing: OTS from RESIDE-$\beta$ SOTS and SOTS-Outdoor SOTS; Deraining: Rain100L Rain100L; Deblurring: GoPro GoPro; Low-light Enhancement: LOL LOL. Details of these datasets are summarized in appendix. For composited degradation, we use CDD11 dataset DBLP:conf/eccv/GuoGLZLH24. For remote sensing AiOIR, we choose MDRS-Landsat DBLP:journals/inffus/LiheYHJXCSZ25. For these two settings, we follow the original split of training and test sets.
@@
Following DBLP:conf/iclr/0001ZKKSK25, we adopt task-specific resampling ratios to address data imbalance across restoration tasks. Specifically, the data expansion ratios of 3, 120, 5, and 200 are applied to denoising, deraining, deblurring, and low-light enhancement, respectively, while dehazing remains unaltered. Training is conducted on 4 NVIDIA A100 40GB GPUs using cropped patches of size $128\times 128$, with random horizontal and vertical flipping for data augmentation.
Table 1: Performance comparison (PSNR/SSIM) under All-in-One (“N+H+R”) setting. Results are partially sourced from Perceive-IR DBLP:journals/corr/abs-2408-15994.
@@ 4.2 All-in-One Restoration Results
Tab.1 shows that EvoIR achieves the best average PSNR (33.00 dB) and SSIM (0.922). Compared to the three latest AiOIR methods: AdaIR DBLP:conf/iclr/0001ZKKSK25, MoCE-IR DBLP:conf/cvpr/ZamfirWMTP0T25 and DFPIR DBLP:conf/cvpr/TianLLLR25, our method improves the average PSNR/SSIM by +0.31 dB/+0.004, +0.23 dB/+0.005 and +0.12 dB/+0.003, respectively. Notably, EvoIR achieves new state-of-the-art results on dehazing (32.08/0.982) and deraining (39.07/0.985), demonstrating the effectiveness of adaptive frequency modulation and evolutionary optimization.
@@
Table 2: Performance comparison (PSNR/SSIM) under All-in-One (“N+H+R+B+L”) setting. Denoising results only reports noise level under $\sigma=25$ following IDR. Results are partially sourced from Perceive-IR DBLP:journals/corr/abs-2408-15994.
@@ Appendix B Code and Models
Code and models for evaluation are released: https://github.com/leonmakise/EvoIR.
@@ C.2 Baselines
All-in-One setting: For “N+H+R” setting, we include the following recent All-in-One models: AirNet AirNet, IDR IDR, ProRes ProRes, PromptIR PromptIR, NDR NDR, Gridformer Gridformer, InstructIR InstructIR, Up-Restorer DBLP:conf/aaai/LiuYL025, Perceive-IR DBLP:journals/corr/abs-2408-15994, AdaIR DBLP:conf/iclr/0001ZKKSK25, MoCE-IR DBLP:conf/cvpr/ZamfirWMTP0T25, DFPIR DBLP:conf/cvpr/TianLLLR25. For “N+H+R+B+L” setting, we also include TAPE TAPE, TransWeather Transweather.
One-by-One setting: We adopt task-specific and general methods tailored to individual degradation types. For Denoising: DnCNN DnCNN, FFDNet FFDNet, ADFNet ADFNet, MIRNet-v2 MIRNet_v2, Restormer Restormer, NAFNet NAFNet; Dehazing: adding DehazeNet DehazeNet, FDGAN FDGAN, DehazeFormer DehazeFormer, FSNet FSNet; Deraining: adding UMR UMR, LPNet LPNet, DRSformer DRSformer; Deblurring: adding DeblurGAN DeblurGAN, Stripformer Stripformer, HI-Diff HI-Diff, MPRNet MPRNet; Low-light enhancement: adding URetinex URetinex, SMG SMG, Retinexformer Retinexformer, MIRNet MIRNet, DiffIR DiffIR. We report All-in-One methods retrained under the One-by-One setting for comparison.
@@ E.2 Visual Comparison under Three Degradations
Due to page limitations, we provide additional visual comparisons for the three-degradation setting (“N+H+R”). As shown in Fig. 6, 7, 8, 9, and 10, we compare our EvoIR against recent state-of-the-art methods, including AdaIR DBLP:conf/iclr/0001ZKKSK25 (ICLR’25), DFPIR DBLP:conf/cvpr/TianLLLR25 (CVPR’25), and MoCE-IR DBLP:conf/cvpr/ZamfirWMTP0T25 (CVPR’25).
@@ E.3 Visual Comparison under Five Degradations
As some methods do not provide visual results under the five-degradation setting (“N+H+R+B+L”), we select InstructIR InstructIR (ECCV’24) and MoCE-IR DBLP:conf/cvpr/ZamfirWMTP0T25 (CVPR’25) for comparison with our EvoIR.